Alaskan Fisheries, Inc., processes salmon for various distributors and it uses the weighted-average method in its process costing system. The company has two processing departments—Cleaning and Packing. Data relating to pounds of salmon processed in the Cleaning Department during July are presented below:

Alaskan Fisheries, Inc., processes salmon for various distributors and it uses the weighted-average method in its process costing system. The company has two processing departments—Cleaning and Packing. Data relating to pounds of salmon processed in the Cleaning Department during July are presented below:

Percent Completed
Pounds of Salmon Materials Labor and Overhead
Work in process inventory, July 1 29,000 100 % 53 %
Work in process inventory, July 31 20,000 100 % 70 %

A total of 460,000 pounds of salmon were started into processing during July. All materials are added at the beginning of processing in the Cleaning Department.

Required:

Compute the Cleaning Department's equivalent units of production for materials and for labor and overhead in the month of July

A total of 460,000 pounds of salmon were started into processing during July. All materials are added at the beginning of processing in the Cleaning Department.
Answer #1

A B с D E F 1 2 Tons completed and transferred out 3 Work in process, July 31 4 Equivalent units of production 5 6 Final answ

A B С D E F Units % Materials EUP-Materials % Conversion 1 EUP- Labor and Overhead =+B2*E2 =+B2*C2 1 0.7 =+B3*C3 =+D2+D3 =+B3